# An Explainable Framework for Particle Swarm Optimization using Landscape Analysis and Machine Learning

### v1

@@ Abstract
Swarm intelligence algorithms have demonstrated remarkable success in solving complex optimization problems across diverse domains. However, their widespread adoption is often hindered by limited transparency in how algorithmic components influence performance. This work presents a multi-faceted investigation of Particle Swarm Optimization (PSO) to further understand the key role of different topologies for better interpretability and explainability. To achieve this objective, we first develop a comprehensive landscape characterization framework using Exploratory Landscape Analysis (ELA) to quantify problem difficulty and identify critical features affecting the optimization performance of PSO. Next, we conduct a rigorous empirical study comparing three fundamental swarm communication architectures - Ring, Star, and Von Neumann topologies - analysing their distinct impacts on exploration-exploitation balance, convergence behaviour, and solution quality and eventually develop an explainable benchmarking framework for PSO, to decode how swarm topologies affects information flow, diversity, and convergence. Based on this, a novel machine learning approach for automated algorithm configuration is introduced for training predictive models on extensive Area over the Convergence Curve (AOCC) data to recommend optimal settings based on problem characteristics. Through systematic experimentation across twenty four benchmark functions in multiple dimensions, we establish practical guidelines for topology selection and parameter configuration. These findings advance the development of more transparent and reliable swarm intelligence systems. The source codes of this work can be accessed at https://github.com/GitNitin02/ioh_pso.
@@ 1 Introduction
Swarm Intelligence (SI) represents a fascinating and powerful field of study that draws inspiration from the collective behaviour observed in decentralized, self-organized natural systems [7]. The complex collective intelligence observed in nature, from the synchronized movements of bird flocks to the optimized foraging of ant colonies, provides the foundational inspiration for sophisticated optimization algorithms. At its core, Swarm Intelligence (SI) is characterized by a population of simple, autonomous agents that follow basic rules and interact locally with their environment and neighbors. Without any centralized guidance, these limited interactions collectively produce sophisticated global patterns and solve complex problems. The approach is fundamentally defined by principles such as decentralization, emergence [7] , adaptation, and positive feedback. Any swarm based optimization algorithm includes three fundamental components: (i) a population that evolves iteratively, (ii) a reproduction process where current population of candidate solutions generates new set of candidate solutions in the next generation, and (iii) a curated information sharing along with an efficient learning mechanism that determines next generation of candidate solutions. These algorithms are inherently scalable, adaptable, and robust, characterized by the simplicity of their individual candidate solutions, making them highly effective for solving a wide array of complex optimization problems [8, 4]. The efficacy of SI algorithms is fundamentally dependent on a critical balance: the exploration of uncharted regions of the search space to discover novel solutions, and the exploitation of known promising areas to refine them. Conversely, exploitation involves refining the search around known promising solutions, leveraging existing information to converge efficiently. Achieving an optimal balance between these two dynamics is paramount for efficiently converging to acceptable solutions while simultaneously avoiding premature convergence to suboptimal local optima. This decentralized architecture is the direct source of the method’s primary strengths: robustness and adaptability [7]. Unlike traditional top-down optimization, the overall solution emerges from the bottom-up interactions of the swarm. This emergent quality makes such systems intrinsically flexible and resilient to dynamic, high-dimensional problem spaces, as the loss of any single agent does not derail the collective process. If one candidate solution fails or a local condition changes, the overall swarm can still adapt and find solutions because its intelligence is distributed, not concentrated in a single vulnerable point. This distributed intelligence makes SI algorithms particularly well-suited for problems where the environment is uncertain, dynamic, or too complex for a single, pre-programmed control mechanism [14]. Among the myriad of SI algorithms, several have gained prominence due to their effectiveness and versatility. Particle Swarm Optimization (PSO) [22] is a prominent algorithm in this category, modeled on the social dynamics of species like birds and fish. In PSO, particles (candidate solutions) traverse a search space, continually adjusting their positions based on their personal best (pbest) known position and the global best (gbest) position discovered by the entire swarm. PSO is particularly well-suited for continuous optimization problems [27], such as the crucial task of parameter tuning [20] in machine learning models. There are several other swarm based optimization algorithms such as artificial bee colony (ABC)[21], ant colony optimization (ACO) [11] and many more but our focus in this article is particle swarm optimization (PSO) which is one of the earliest swarm based optimization algorithms. Parameters play a fundamental role in the effective performance of swarm based optimization algorithms, profoundly influencing a meta-heuristic’s efficiency in solving a given decision problem [2]. The primary function of these parameters is to carefully balance the convergence and diversity of the search process, which directly translates to managing the trade-off between exploration (searching new areas of the solution space) and exploitation (refining the search around currently identified promising solutions) [43]. Suboptimal selection of these control parameters can lead to significant issues, such as premature convergence, where the swarm settles on a suboptimal solution too quickly, and an increase in computational time. The quality of the solution, the amount of computational power needed, the degree of diversity and exploration attained, and the general leadership dynamics within the swarm are all directly impacted by parameter settings [8]. Putting our focus back on PSO, its performance is notably sensitive to its control parameters, particularly the inertia weight $(w)$ and the acceleration coefficients $(c_{1},c_{2})$ [43].Throughout the search process, dynamically balancing exploration and exploitation requires adaptive strategies for these parameters. Our comprehension of PSO parameters is further improved by the following points:
  1. 1.
@@
1.1 Need of Explainabilty in PSO
@@
Despite its demonstrated efficacy and widespread adoption, PSO, akin to many advanced AI algorithms, it (PSO) frequently functions as a “black box,” [53] obscuring its internal decision-making processes. This inherent lack of clarity poses significant challenges for building trust and verifying the reliability of its outcomes, particularly when applied in sensitive or critical domains like healthcare and autonomous systems. The ambiguous nature of PSO’s configurations and hyperparameters can directly contribute to a perceived lack of reliability in the solutions generated by the algorithm. The stochastic nature of particle movements and the emergent, non-linear interactions within the swarm contribute significantly to this opacity, making it difficult for human users to understand why a specific optimal solution was found or how the swarm navigated the complex search space to converge. This apparent contradiction between PSO’s algorithmic simplicity and its behavioural opacity, reveals a deeper truth: PSO’s algorithmic rules are simple to implement, yet its collective behavior is highly complex and inscrutable. This obscurity does not arise from intricate internal structures but from the swarm’s emergent “intelligence”—the intricate interplay of local and global information sharing across a high-dimensional space. The challenge for XAI is thus that the “black box” nature of such metaheuristics stems from the non-linear dynamics of many simple agents, which confounds a clear tracing of the causal pathway to a solution. Therefore, XAI methods for PSO need to focus on analyzing collective behavior and parameter sensitivity rather than just decomposing complex architectures. In response to this challenge, the fields of Explainable Artificial Intelligence (XAI) and Interpretable Artificial Intelligence (IAI) [12] have rapidly gained prominence. Their primary objective is to develop methodologies that render AI models inherently understandable, thereby clarifying their decision-making processes and nurturing greater confidence in their results.
@@
Historically, metaheuristic optimization algorithms like PSO were primarily evaluated based on their efficiency, convergence speed, and the quality of the solutions they found. However, recent work [16] on analysing the impact of swarm topologies provides an early work in the direction of “enhancing explainability and trustworthiness” in PSO. This work is an extension of the work carried out in [16]. This signals a significant maturation within the field of computational intelligence. It indicates that achieving optimal solutions is no longer the sole criterion for algorithm success. Instead, there is a growing recognition that for AI systems to be responsibly deployed and widely adopted, especially in critical applications, they must also be transparent and understandable [3]. This reflects a broader ethical and practical demand for trustworthy AI. This paradigm shift implies a pressing need for the development of new evaluation metrics that extend beyond traditional optimization performance (e.g., solution quality, computational cost) to quantitatively assess explainability. It also suggests that future research in PSO will increasingly integrate XAI considerations from the very initial design phases of new algorithms, rather than merely applying post-hoc explanations [45].
1.1.1 Analyzing the Impact of Communication Topologies in PSO
@@
The effectiveness and search behavior of PSO are profoundly influenced by its communication topology, which defines the structural pattern through which particles share information within the swarm. Common topologies include Ring, [29, 47] Star [37, 38], and Von Neumann [51, 34]. Research investigates how these distinct topologies impact critical aspects of PSO’s operation, such as information flow, diversity maintenance, convergence speed, and the delicate balance between exploration (searching new regions) and exploitation (refining known good regions). By systematically analyzing these influences, it is possible to enhance the inherent interpretability of PSO’s decision-making processes and derive practical guidelines for selecting the most suitable topology for specific optimization tasks. The explicit detailing of how different communication topologies directly affect PSO’s convergence, diversity, and exploration-exploitation balance provides a powerful form of intrinsic interpretability. This approach is powerful because it moves beyond treating PSO as an indivisible black box. Instead, it focuses on understanding and designing its core components to be inherently transparent. This suggests that modifying the algorithm’s internal architecture and interaction rules can directly contribute to its interpretability, offering a promising avenue for developing more transparent and trustworthy metaheuristics, in future.
@@ 1.1.2 Explainable Benchmarking Frameworks for PSO
1.1.3 Understanding Benchmark Problems
@@
  1. 1.
    We utilize Explainable Landscape Analysis (ELA) to compute and visualize key features (e.g., modality, smoothness, variable interactions) of the Black Box Optimization Benchmark (BBOB) problems, creating an interpretable “fingerprint” for each function’s structure.
  2. 2.
    We construct a comprehensive PSO hyperparameter configuration space containing all necessary hyperparameters for Star, Ring, and Von Neumann topologies.
  3. 3.
    We integrate the configurable PSO into the IOHxplainer framework to generate Area over the Convergence Curve (AOCC) performance data, enabling an analysis of hyperparameter importance and it’s contribution.
  4. 4.
    We train a Decision Tree and Random Forest classifier on the run data to predict performance classes. The resulting tree is visualized to show the feature split at each node, including the distribution of the target class (e.g., high/low performance) for the feature’s values, enabling the identification of critical hyperparameter thresholds and optimal decision rules.
This work is an extension of our own work on explainability of PSO topologies [16]. This paper is organized as follows: Section 2 covers ELA-based visualization for benchmarking. Section 3 proposes a novel explainability framework for PSO. Section 4 presents the experimental setup and results, and Section 5 analyzes the impact of hyperparameters on performance. Section 6 describes our method for data-driven, interpretable configuration learning. Finally, Section 7 concludes and suggests future work.
@@ 2 ELA-based visualization for benchmarking
Building on these cell mapping foundations, more advanced constructs like barrier trees [25] can be generated. In this representation, local optima (valleys) are depicted as leaves (filled circles), while ridges connecting adjacent valleys are represented by branching nodes (non-filled diamonds), providing a hierarchical model of the fitness landscape’s structure.Visualization of the cell mapping plots and barrier trees plot are shown in Tables 1-3. Similarly, as detailed in Table 4, Information Content of Fitness Sequences (ICoFiS) measures properties such as smoothness, ruggedness, and neutrality by analyzing the entropy in fitness sequences generated from random walks with variable step sizes. This approach effectively adapts discrete information-theoretic measures for continuous optimization domains, with the metrics in Table 4 serving as a direct output of this analysis. The comprehensive nature of this approach is summarized in Tables 5-6, which details feature sets such as those computed by calculate_ela_meta, calculate_ela_distribution, calculate_nbc, calculate_dispersion, and calculate_information_content used to capture the global structure, value distribution, clustering behavior, dispersion, and informational complexity of a selected BBOB function’s landscape. Section 6 details a classification methodology where an expanded suite of ELA features (building on Tables 5 and 6) is used to predict categorical performance of the PSO algorithm, as measured by its AOCC.

### v2

# Explainable Framework for Swarm Intelligence Based on Fitness Landscape Features and Machine Learning

@@ Abstract
Swarm based optimization algorithms have demonstrated remarkable success in solving complex optimization problems. However, their widespread adoption remains sceptical due to limited transparency in how different algorithmic components influence the overall performance of the algorithm. This work presents a multi-faceted interpretability related investigations of one of the popular swarm algorithms, Particle Swarm Optimization. Through this work, we provide a framework that makes the role of different topologies and parameters in PSO interpretable and explainable using novel machine learning approach. We first developed a comprehensive landscape characterization framework using Exploratory Landscape Analysis to quantify problem difficulty and identify critical features in the problem that affects the optimization performance of PSO. Secondly, we rigorously compare three topologies - Ring, Star, and Von Neumann analyzing their distinct impacts on exploration-exploitation balance, convergence behavior, and solution quality and eventually develop an explainable benchmarking framework for PSO. The work successfully decodes how swarm topologies affect information flow, diversity, and convergence. Through systematic experimentation across 24 benchmark functions in multiple dimensions, we establish practical guidelines for topology selection and parameter configuration. These findings uncover the black-box nature of PSO, providing more transparency and interpretability to swarm intelligence systems. The source code is available at https://github.com/GitNitin02/ioh_pso.
@@ 1 Introduction
Swarm Intelligence (SI) represents a fascinating and powerful field of study that draws inspiration from the collective behaviour observed in decentralized, self-organized natural systems [7]. The complex collective intelligence observed in nature, from the synchronized movements of bird flocks to the optimized foraging of ant colonies, provides the foundational inspiration for sophisticated optimization algorithms. At its core, Swarm Intelligence (SI) is characterized by a population of simple, autonomous agents that follow basic rules and interact locally with their environment and neighbors. Without any centralized guidance, these limited interactions collectively produce sophisticated global patterns and solve complex problems. The approach is fundamentally defined by principles such as decentralization, emergence [7] , adaptation, and positive feedback. Any swarm based optimization algorithm includes three fundamental components: (i) a population that evolves iteratively, (ii) a reproduction process where current population of candidate solutions generates new set of candidate solutions in the next generation, and (iii) a curated information sharing along with an efficient learning mechanism that determines next generation of candidate solutions. These algorithms are inherently scalable, adaptable, and robust, characterized by the simplicity of their individual candidate solutions, making them highly effective for solving a wide array of complex optimization problems [8, 4, 42]. The efficacy of SI algorithms is fundamentally dependent on a critical balance: the exploration of uncharted regions of the search space to discover novel solutions, and the exploitation of known promising areas to refine them. Conversely, exploitation involves refining the search around known promising solutions, leveraging existing information to converge efficiently. Achieving an optimal balance between these two dynamics is paramount for efficiently converging to acceptable solutions while simultaneously avoiding premature convergence to suboptimal local optima. This decentralized architecture is the direct source of the method’s primary strengths: robustness and adaptability [7]. Unlike traditional top-down optimization, the overall solution emerges from the bottom-up interactions of the swarm. This emergent quality makes such systems intrinsically flexible and resilient to dynamic, high-dimensional problem spaces, as the loss of any single agent does not derail the collective process. If one candidate solution fails or a local condition changes, the overall swarm can still adapt and find solutions because its intelligence is distributed, not concentrated in a single vulnerable point. This distributed intelligence makes SI algorithms particularly well-suited for problems where the environment is uncertain, dynamic, or too complex for a single, pre-programmed control mechanism [15]. Among the myriad of SI algorithms, several have gained prominence due to their effectiveness and versatility. Particle Swarm Optimization (PSO) [23] is a prominent algorithm in this category, modeled on the social dynamics of species like birds and fish. In PSO, particles (candidate solutions) traverse a search space, continually adjusting their positions based on their personal best (pbest) known position and the global best (gbest) position discovered by the entire swarm. PSO is particularly well-suited for continuous optimization problems [29], such as the crucial task of parameter tuning [21] in machine learning models. There are several other swarm based optimization algorithms such as artificial bee colony (ABC)[22], ant colony optimization (ACO) [12] and many more but our focus in this article is particle swarm optimization (PSO) which is one of the earliest swarm based optimization algorithms. Parameters play a fundamental role in the effective performance of swarm based optimization algorithms, profoundly influencing a meta-heuristic’s efficiency in solving a given decision problem [2]. The primary function of these parameters is to carefully balance the convergence and diversity of the search process, which directly translates to managing the trade-off between exploration (searching new areas of the solution space) and exploitation (refining the search around currently identified promising solutions) [48]. Suboptimal selection of these control parameters can lead to significant issues, such as premature convergence, where the swarm settles on a suboptimal solution too quickly, and an increase in computational time. The quality of the solution, the amount of computational power needed, the degree of diversity and exploration attained, and the general leadership dynamics within the swarm are all directly impacted by parameter settings [8]. Putting our focus back on PSO, its performance is notably sensitive to its control parameters, particularly the inertia weight $(w)$ and the acceleration coefficients $(c_{1},c_{2})$ [48].Throughout the search process, dynamically balancing exploration and exploitation requires adaptive strategies for these parameters. Our comprehension of PSO parameters is further improved by the following points:
  1. i.
@@
1.1 Need of Explainabilty in PSO and Current Status of Research
@@
Despite its demonstrated efficacy and widespread adoption, PSO, akin to many advanced AI algorithms, it (PSO) frequently functions as a “black box,” [59] obscuring its internal decision-making processes. This inherent lack of clarity poses significant challenges for building trust and verifying the reliability of its outcomes, particularly when applied in sensitive or critical domains like healthcare and autonomous systems. Current literature reveals a multi-faceted approach to interpreting these complex, self-organizing systems. A prominent strategy involves post-hoc explanation methods, where surrogate models—such as decision trees or linear models—are trained on the input-output data generated by the swarm to provide a global, simplified summary of solution features [30]. Complementing this, social network analysis (SNA) is used to model the dynamic communication topology of a swarm, quantifying the influence of individual agents and tracing the flow of information that leads to convergence [24]. For a more intrinsic understanding, fitness landscape analysis (FLA) characterizes the problem’s topology to explain algorithmic performance, such as justifying convergence stalls by identifying multimodal ruggedness [38]. Furthermore, visual analytics and dimensionality reduction techniques are leveraged to create interactive dashboards and project high-dimensional swarm trajectories into 2D or 3D spaces, allowing users to visually track exploration and exploitation behavior over time [61]. Despite these advances, challenges remain in standardizing evaluation metrics for explanations and developing real-time, human-in-the-loop XAI frameworks to fully bridge the gap between powerful optimization and user trust. The ambiguous nature of PSO’s configurations and hyperparameters can directly contribute to a perceived lack of reliability in the solutions generated by the algorithm. The stochastic nature of particle movements and the emergent, non-linear interactions within the swarm contribute significantly to this opacity, making it difficult for human users to understand why a specific optimal solution was found or how the swarm navigated the complex search space to converge. This apparent contradiction between PSO’s algorithmic simplicity and its behavioural opacity, reveals a deeper truth: PSO’s algorithmic rules are simple to implement, yet its collective behavior is highly complex and inscrutable. This obscurity does not arise from intricate internal structures but from the swarm’s emergent “intelligence”—the intricate interplay of local and global information sharing across a high-dimensional space. The challenge for XAI is thus that the “black box” nature of such metaheuristics stems from the non-linear dynamics of many simple agents, which confounds a clear tracing of the causal pathway to a solution. Therefore, XAI methods for PSO need to focus on analyzing collective behavior and parameter sensitivity rather than just decomposing complex architectures. In response to this challenge, the fields of Explainable Artificial Intelligence (XAI) and Interpretable Artificial Intelligence (IAI) [13] have rapidly gained prominence. Their primary objective is to develop methodologies that render AI models inherently understandable, thereby clarifying their decision-making processes and nurturing greater confidence in their results.
@@
Historically, metaheuristic optimization algorithms like PSO were primarily evaluated based on their efficiency, convergence speed, and the quality of the solutions they found. However, recent work [17] on analysing the impact of swarm topologies provides an early work in the direction of “enhancing explainability and trustworthiness” in PSO. This work is an extension of the work carried out in [17]. This signals a significant maturation within the field of computational intelligence. It indicates that achieving optimal solutions is no longer the sole criterion for algorithm success. Instead, there is a growing recognition that for AI systems to be responsibly deployed and widely adopted, especially in critical applications, they must also be transparent and understandable [3]. This reflects a broader ethical and practical demand for trustworthy AI. This paradigm shift implies a pressing need for the development of new evaluation metrics that extend beyond traditional optimization performance (e.g., solution quality, computational cost) to quantitatively assess explainability. It also suggests that future research in PSO will increasingly integrate XAI considerations from the very initial design phases of new algorithms, rather than merely applying post-hoc explanations [50]. The follwoing subsections further throws light on some key factors which have the highest impact in the interpretability and explainability of PSO.
1.1.1 Impact of Communication Topologies in PSO
@@
The effectiveness and search behavior of PSO are profoundly influenced by its communication topology, which defines the structural pattern through which particles share information within the swarm. Common topologies include Ring, [32, 52] Star [41, 43], and Von Neumann [56, 37]. Research investigates how these distinct topologies impact critical aspects of PSO’s operation, such as information flow, diversity maintenance, convergence speed, and the delicate balance between exploration (searching new regions) and exploitation (refining known good regions) [60]. By systematically analyzing these influences, it is possible to enhance the inherent interpretability of PSO’s decision-making processes and derive practical guidelines for selecting the most suitable topology for specific optimization tasks. The explicit detailing of how different communication topologies directly affect PSO’s convergence, diversity, and exploration-exploitation balance provides a powerful form of intrinsic interpretability. This approach is powerful because it moves beyond treating PSO as an indivisible black box. Instead, it focuses on understanding and designing its core components to be inherently transparent. This suggests that modifying the algorithm’s internal architecture and interaction rules can directly contribute to its interpretability, offering a promising avenue for developing more transparent and trustworthy metaheuristics, in future.
@@ 1.1.2 Explainable Benchmarking Frameworks for PSO
1.1.3 Deeper Understanding Benchmark Problems
@@
1.2 Highlights of Identified Research Gaps
Despite considerable progress in understanding Particle Swarm Optimization (PSO) and its communication topologies, significant research gaps persist in making PSO explainable, generalizable, and trustworthy. Existing studies largely rely on empirical performance benchmarking without sufficiently explaining the causal mechanisms driving PSO’s behavior under varied landscapes. The lack of a unified explainability framework that connects PSO’s hyperparameters, topological structures, and problem-specific landscape features limits our capacity to derive interpretable insights.
The integration of IOHxplainer and Exploratory Landscape Analysis (ELA) offers a significant advancement towards bridging the research gaps identified in explainable and trustworthy PSO systems. These two frameworks complement each other by providing both algorithm-centric and problem-centric perspectives on optimization behaviour.
The proposed research work contribute to filling the following key research gaps by:
  1. i.
    Establishing quantitative and interpretable mappings between PSO parameters, topologies, and performance.
  2. ii.
    Providing reproducible, explainable benchmarking methodologies beyond mere empirical comparison.
  3. iii.
    Enabling visualization-driven insights into swarm dynamics and problem structures.
  4. iv.
    Laying the foundation for developing adaptive and transparent optimization algorithms.
  5. v.
    Promoting generalization and scalability through data-driven interpretability across multiple problem classes.
1.3 Summary of Contributions
This work makes several significant contributions towards enhancing the explainability, interpretability, and trustworthiness of the PSO framework by integrating insights from machine learning and landscape analysis. The work introduces a novel explainable modeling pipeline that systematically connects PSO’s algorithmic parameters, communication topologies, and problem landscape characteristics to its performance outcomes. By combining Exploratory Landscape Analysis (ELA) with the IOHxplainer framework, the work provides both problem-level and algorithm-level interpretability, thereby addressing the long-standing “black-box” nature of PSO.
The key contributions of the manuscript can be summarized as follows:
  1. i.
    Proposed a novel explainable PSO framework integrating Exploratory Landscape Analysis (ELA) and IOHxplainer for interpretable optimization.
  2. ii.
    Linked PSO parameters, communication topologies, and landscape features to algorithmic performance for deeper causal understanding.
  3. iii.
    Analyzed three topologies Star, Ring, and Von Neumann to explain their impact on exploration exploitation balance and convergence.
  4. iv.
    Utilized IOHxplainer with SHAP-based analysis for explainable benchmarking and hyperparameter impact quantification.
  5. v.
    Employed ELA for problem characterization, identifying structural features influencing optimization difficulty.
  6. vi.
    Developed a framework for explainability of PSO thorough decision tree.
  7. vii.
    Conducted extensive experiments across 24 BBOB benchmark functions to validate the framework’s generalizability.
  8. viii.
    Advanced the concept of trustworthy and transparent swarm intelligence through interpretable algorithmic design.
@@ 1.3 Summary of Contributions
  1. i.
    The work employs Exploratory Landscape Analysis (ELA) to compute and visualize critical landscape features—such as modality, smoothness, and variable interactions of the Black-Box Optimization Benchmark (BBOB) functions, thereby generating an interpretable fingerprint that characterizes each problem’s structural complexity.
  2. ii.
    A comprehensive PSO hyperparameter configuration space is formulated, encompassing all essential control parameters corresponding to the Star, Ring, and Von Neumann communication topologies, to facilitate systematic and comparative analysis.
  3. iii.
    The configurable PSO framework is integrated within the IOHxplainer environment to produce Area Over the Convergence Curve (AOCC) performance data, enabling a rigorous evaluation of hyperparameter significance and their respective contributions to algorithmic performance.
  4. iv.
    Decision Tree and Random Forest classifiers are trained on the experimental run data to predict distinct performance classes. The resulting interpretable decision trees visualize feature splits and corresponding class distributions, thereby revealing critical hyperparameter thresholds and optimal decision rules that govern PSO behavior.
This work extends our earlier work on the explainability of PSO communication topologies [17], further enhancing interpretability through the integration of ELA and IOHxplainer-based analysis. The remainder of this paper is organized as follows: Section 2 presents the ELA-based visualization framework employed for benchmarking. Section 3 introduces the proposed explainability framework for PSO, followed by Section 4, which details the experimental setup and corresponding results. The influence of hyperparameters on algorithmic performance is examined in Section 5, while Section 6 outlines the data-driven, interpretable configuration learning methodology. Finally, Section 7 provides the concluding remarks and potential directions for future research.
@@ 2 ELA-based visualization for benchmarking
Building on these cell mapping foundations, more advanced constructs like barrier trees [27] can be generated. In this representation, local optima (valleys) are depicted as leaves (filled circles), while ridges connecting adjacent valleys are represented by branching nodes (non-filled diamonds), providing a hierarchical model of the fitness landscape’s structure.Visualization of the cell mapping plots and barrier trees plot are shown in Tables 1-3. Similarly, as detailed in Table 4, Information Content of Fitness Sequences (ICoFiS) measures properties such as smoothness, ruggedness, and neutrality by analyzing the entropy in fitness sequences generated from random walks with variable step sizes. This approach effectively adapts discrete information-theoretic measures for continuous optimization domains, with the metrics in Table 4 serving as a direct output of this analysis. The comprehensive nature of this approach is summarized in Tables 5-6, which details feature sets such as those computed by calculate_ela_meta, calculate_ela_distribution, calculate_nbc, calculate_dispersion, and calculate_information_content used to capture the global structure, value distribution, clustering behaviour, dispersion, and informational complexity of a selected BBOB function’s landscape. Section 6 details a classification methodology where an expanded suite of ELA features (building on Tables 5 and 6) is used to predict categorical performance of the PSO algorithm, as measured by its AOCC. The remaining set of ELA features, used for the classification task, has been made publicly available on GitHub.